C(10,5) x C(15,8)
__________________
C(12,11)

Answer by checkley75(3666)

C(10,5) x C(15,8)
__________________
C(12,11)
10C5=252
15C8=6,435
12C11=12
252*6,435/12=1,621,620/12=135,135 ANSWER.